Determine the smallest size of hole that can be punched in 12 mm thick MS 

plate having ultimate shear stress of 390 N/mm2 , if permissible crushing 

stress for the punch material is 2.4 kN/mm2.


1
Expert's answer
2021-01-06T06:04:12-0500

Ultimate shear stress="\tau_u=" 390 N/mm2, thickness= 12mm,


Permissible crushing stress=2.4 kN/mm2


Axial shear force= "(\frac{\pi d^2}{4})\tau= F" ,In case of crushing


"F=\pi dt \times \sigma_c" as we know that in case of small size both force will be same


"\pi dt \sigma_c=\frac{\pi d^2}{4}\times \tau_u"


12 "\times 2400= \frac{d}{4}\times 390"


d=295.38 mm